Can someone please help me with this type of problem?(-3/7xyz)(x^2y^3)(7xyz^2).
Ty

Answer by edjones(8007)

(-3/7xyz)(x^2y^3)(7xyz^2)
Let's take the coefficients 1st:
-3/7*7= -3
Now the variables one at a time. Add exponents when multiplying variables.
x*x^2*x= x^4
y*y^3*y= y^5
z*z^2= z^3
.
-3x^4y^5z^3 Answer.
.
Ed